# version 1.0

### v6

@@ 11.1 Spectral filtering in One Dimension
The one-dimensional systems which we consider next are scalar systems, with $d_{u}=d_{y}=d_{x}=1$. For simplicity, we assume that the initial state is zero $\mathbf{x}_{0}=0$. Recall the set of optimal predictors, the class $\Pi_{\star}$, which for scalar systems can written and simplified to,
$\mathbf{y}_{t}$ $=\sum_{i=1}^{t}CA^{i-1}B\mathbf{u}_{t-i}=\beta\sum_{i=1}^{t}\alpha^{i-1}\mathbf{u}_{t-i},$
Where $\beta=CB\in{\mathbb{R}}$. Thus, we can write the difference between two consecutive iterates as
$\mathbf{y}_{t}-\mathbf{y}_{t-1}$ $=\beta(\sum_{i=1}^{t}\alpha^{i-1}\mathbf{u}_{t-i}-\sum_{i=1}^{t-1}\alpha^{i-1}\mathbf{u}_{t-1-i})$
$=\beta(\mathbf{u}_{t-1}+\sum_{i=1:t-1}(\alpha^{i}-\alpha^{i-1})\mathbf{u}_{t-1-i})$
$=\beta\mathbf{u}_{t-1}+\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1}),$ (11.1)
which is a geometric series involving $\alpha$ multiplying the past controls. We used the operator notation to denote the operator
$\mu_{\alpha}(\mathbf{u}_{t-1:1})=\sum_{i=1:t-1}(\alpha-1)\alpha^{i-1}\mathbf{u}_{t-1-i}=\mu_{\alpha}^{\top}\tilde{\mathbf{u}}_{t-1}.$ (11.2)
Above we used the vector notation for
$\mu_{\alpha}=(\alpha-1)\begin{bmatrix}1&\alpha&\alpha^{2}&\ldots&\alpha^{T-1}\end{bmatrix}\in{\mathbb{R}}^{T},$
@@
$\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}.$
@@
The matrix $Z_{T}$ is symmetric and is independent of $\alpha$. Let $\phi_{1},...,\phi_{k},...\in{\mathbb{R}}^{T}$ be the eigenvectors of $Z_{T}$ ordered by the magnitude of their corresponding eigenvalues. Our basis for representing $\mu_{\alpha}$ is spanned by the eigenvectors of $Z_{T}$, and we proceed to study their properties.
@@ 11.1.1 The Magic of Hankel Matrices
The matrix $Z_{T}$ is a special case of a Hankel Matrix, which is a square Hermitian matrix with each ascending skew-diagonal from left to right having the same value. An example of a Hankel matrix is the following,
$W_{T}$ $=\int_{0}^{1}\begin{bmatrix}1&\alpha&\alpha^{2}&&\\ \alpha&\alpha^{2}&\alpha^{3}&&\\ \alpha^{2}&\alpha^{3}&\alpha^{4}&&\\ &&&...&\\ &&&&\alpha^{2T-2}\end{bmatrix}d\alpha=\begin{bmatrix}1&\frac{1}{2}&\frac{1}{3}&&\\ \frac{1}{2}&\frac{1}{3}&\frac{1}{4}&&\\ \frac{1}{3}&\frac{1}{4}&\frac{1}{5}&&\\ &&&...&\\ &&&&\frac{1}{2T-1}\end{bmatrix}$.
This matrix is closely related to the $Z_{T}$ matrix we have introduced before, but not equal to it.
@@
Let $\sigma_{k}$ be the $k$-th largest singular value of a Hankel matrix $M\in{\mathbb{R}}^{T\times T}$, then there exists constants $c_{1},c_{2}$ such that :
$\sigma_{k}(M)\leq c_{1}\cdot e^{-\frac{c_{2}k}{\log T}}$
@@ 11.2 Spectral Predictors
Define the class of spectral linear predictors of order $k$ recursively as
@@
Here the operator $\phi_{i}(\mathbf{u}_{t-1})$ takes the inner product of the $i$’th eigenvector of $Z_{T}$ with $\tilde{\mathbf{u}}_{t-1}$, where $\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}$.
Notice that as $h\mapsto T$, this class is equivalent to the class of all linear predictors on the inputs $\prod^{LIN}$. The reader may wonder at this point: Since a linear transformation composed with another linear transform remains linear, what is the benefit here?
The answer is that linear transformations are equivalent when we have full-rank transforms. In our case, however, we consider low-rank approximations to the transform, and the approximation properties of different subspaces can differ drastically! This is the case in our setting: the spectral basis arising from the matrix $Z_{T}$ gives excellent approximations to the class of (standard) linear predictors, as we formally prove next. The following lemma is analogous to Lemma 9.4, with the significant difference that the number of parameters $h$ does not depend on the stabilizability parameter of the system $\gamma$ at all!
@@
Suppose that $\mathbf{x}_{0}=0$, then the class $\prod_{h}^{\lambda}$ for $h=O(\log^{2}\frac{T}{\varepsilon})$, $\varepsilon$-approximates the class $\prod_{\star}$.
The proof of this lemma relies on the following crucial property of the spectral basis, whose proof we defer to later.
Lemma 11.4.
Let $\{\phi_{j}\}$ be the eigenvectors of Z. Then there exist constants $c_{1},c_{2}>0$ such that for all $j\in[T]$, $\alpha\in[0,1]$, we have
$|\phi_{j}^{\top}\mu_{\alpha}|\leq c_{1}e^{-\frac{c_{2}j}{\log T}}.$
Using this property, we can prove Lemma 11.3 by simply shifting to a representation on the spectral basis as follows.
@@
Consider any sequence and any predictor from $\prod_{\star}$, say $\pi_{1}=\pi_{A,B,C,0}$. Following equation (11.1), we can write the optimal predictor in the spectral basis:
$\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\beta\mathbf{u}_{t}$ $=\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1})$
$=\beta\cdot\mu_{\alpha}^{\top}\tilde{\mathbf{u}}_{t-1}$ (11.2)
$=\beta\cdot\mu_{\alpha}^{\top}(\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top})\tilde{\mathbf{u}}_{t-1}$ $\mbox{basis property }\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top}=I$
$=\sum_{i=1}^{T}M_{i}^{\star}\cdot\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}.$ $M_{i}^{\star}=\beta\phi_{i}^{\top}\mu_{\alpha}$
By Lemma 11.4 we have that
$M_{i}^{\star}=\beta\phi_{i}^{\top}\mu_{\alpha}=O(e^{-\frac{i}{\log T}}).$
We proceed by an inductive claim on $\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|$. Consider the predictor $\pi_{2}\in\prod_{h}^{\lambda}$ for the same system matrices $M^{\star}$. Then we have
$\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|-\|\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{2}}\|$
$\leq\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}+\mathbf{y}_{t-1}^{\pi_{2}}\|$ -$\triangle\mbox{-inequality}$
$=\left\|\sum_{i=1}^{T}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}-\sum_{i=1}^{h}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}\right\|$
$=\left\|\sum_{i=h+1}^{T}M^{\star}_{i}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-1}\right\|$
$\leq\sum_{i=h+1}^{T}\|M_{i}^{\star}\||\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-1}|$
$\leq\sum_{i=h+1}^{T}\|M_{i}^{\star}\|\|\phi_{i}\|_{2}\|\tilde{\mathbf{u}}_{t-1}\|_{2}$ Cauchy-Schwartz
$\leq\sqrt{T}\cdot\sum_{i=h+1}^{T}|M_{i}^{\star}|$ $|\mathbf{u}_{t}|\leq 1,\|\phi_{i}\|_{2}\leq 1$
$\leq\sqrt{T}\times O\left(\int_{h}^{\infty}e^{-\frac{i}{\log T}}di\right)$ Lemma 11.4
$\leq\frac{\varepsilon}{T}.$ choice of $h$
Thus, since same argument exactly applies to $\|\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{2}}\|$, we have that
$\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|\leq\sum_{t=1}^{T}\frac{\varepsilon}{T}\leq\varepsilon.$
It remains to prove Lemma 11.4, which is the main technical lemma of this section.
Lemma 11.4.
Consider the scalar function $g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}$ over the interval $[0,1]$. First, notice that by definition of $\phi_{j}$ as the eigenvectors of $Z_{T}$, and $\sigma_{j}$ as the corresponding eigenvalues, we have
$\int_{\alpha=0}^{1}g(\alpha)d\alpha=\int_{\alpha=0}^{1}\big(\phi_{j}^{\top}\mu_{\alpha}\big)^{2}d\alpha=\int_{\alpha=0}^{1}\phi_{j}^{\intercal}\mu_{\alpha}\mu_{\alpha}^{\intercal}\phi_{j}d\alpha=\phi_{j}^{\intercal}Z_{T}\phi_{j}=\sigma_{j}.$
On the other hand, we claim that the function $g(\alpha)$ is L-Lipschitz for a constant value of $L$. To see this, first notice that
$\|\mu_{\alpha}\|_{2}^{2}=\sum_{i=0}^{T-1}(1-\alpha)^{2}\alpha^{2i}\leq\frac{(1-\alpha)^{2}}{1-\alpha^{2}}=\frac{1-\alpha}{1+\alpha}\leq 1-\alpha,$
implying that $g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}\leq\|\phi_{j}\|^{2}\|\mu_{\alpha}\|^{2}\leq 1-\alpha$. Next, we have
$\left\|\frac{\partial}{\partial\alpha}\mu_{\alpha}\right\|_{2}^{2}$ $=\sum_{t=0}^{T-1}((1-\alpha)i\alpha^{i-1}-\alpha^{i})^{2}$
$\leq 2(1-\alpha)^{2}\sum_{t=0}^{T-1}i^{2}\alpha^{2i-2}+2\sum_{t=1}^{T}\alpha^{2i}$ $(a+b)^{2}\leq 2(a^{2}+b^{2})$
$\leq 2(1-\alpha)^{2}\left(\frac{1}{(1-\alpha^{2})^{2}}+\frac{\alpha^{2}}{(1-\alpha^{2})^{3}}\right)+\frac{2}{1-\alpha^{2}}$
$\leq\frac{6}{1-\alpha},$
where we use the identity $\sum_{i=0}^{\infty}i^{2}\beta^{i-1}=\frac{1}{(1-\beta)^{2}}+\frac{\beta}{(1-\beta)^{3}}$ with the substitution $\beta=\alpha^{2}$. Using this, we obtain an upper bound on the Lipschitz constant of $g(\alpha)$ as
$|g^{\prime}{(\alpha)}|$ $=\left\lvert\frac{\partial}{\partial{\alpha}}(\mu_{\alpha}^{\top}\phi_{j})^{2}\right\rvert=2|\mu_{\alpha}^{\top}\phi_{j}|\left\lvert\frac{\partial}{\partial{\alpha}}(\mu_{\alpha}^{\top}\phi_{j})\right\rvert$
$\leq 2\|\mu_{\alpha}\|_{2}\|\phi_{j}\|_{2}\left\lvert\frac{\partial}{\partial{\alpha}}(\mu_{\alpha}^{\top}\phi_{j})\right\rvert$ Cauchy-Schwarz
$\leq 2\|\mu_{\alpha}\|_{2}\|\phi_{j}\|^{2}_{2}\left\|\frac{\partial}{\partial{\alpha}}\mu_{\alpha}\right\|_{2}$ Cauchy-Schwarz
$\leq 2\sqrt{1-\alpha}\sqrt{\frac{6}{1-\alpha}}\leq 6.$ $\|\phi_{i}\|_{2}\leq 1$
Let $g_{\max}=\max_{\alpha}g(\alpha)$ be the maximal value that $g$ attains from zero to one. Any non-negative L-Lipschitz function that reaches $g_{\max}$ and has the area of $A$, satisfies:
$A\geq\frac{g_{\max}^{2}}{2L}\ \ \Rightarrow\ \ g_{\max}^{2}\leq{2LA},$
since the minimal area is realized by a triangle.
Thus, the maximum value of $g(\alpha)$ is bounded by, using Theorem 11.1,
$g_{\max}=\max_{\alpha}\left\{(\phi_{j}^{\top}\mu_{\alpha})^{2}\right\}$ $\leq 2\cdot 6\cdot\int_{\alpha}(\phi_{j}^{\top}\mu_{\alpha})^{2}d\alpha$
$=12\cdot\phi_{j}^{\top}\left(\int_{\alpha}\mu_{\alpha}\mu_{\alpha}^{\top}d\alpha\right)\phi_{j}$
$=12\sigma_{j}$
$\leq 12c_{1}e^{-\frac{c_{2}j}{\log T}}$ Theorem 11.1

### v8

@@ 11.1 Spectral filtering in One Dimension
The one-dimensional systems which we consider next are scalar systems, with $d_{u}=d_{y}=d_{x}=1$. For simplicity, we assume that the initial state is zero $\mathbf{x}_{0}=0$. Recall the set of optimal predictors, the class $\Pi_{\star}$, which for scalar systems can be written and simplified to,
$\mathbf{y}_{t}$ $=\sum_{i=1}^{t-1}CA^{i-1}B\mathbf{u}_{t-i}=\beta\sum_{i=1}^{t-1}\alpha^{i-1}\mathbf{u}_{t-i}=\beta\cdot\tilde{\mathbf{u}}_{t}^{\top}\mu_{\alpha},$ (11.1)
where $\beta=CB\in{\mathbb{R}}$ is a scalar and we used the vector notation for
$\mu_{\alpha}=\begin{bmatrix}1&\alpha&\alpha^{2}&\ldots&\alpha^{T-1}\end{bmatrix}^{\top}\in{\mathbb{R}}^{T},$
@@
$\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0&\ldots&0\end{bmatrix}^{\top}\in{\mathbb{R}}^{T}.$
@@
The matrix $Z_{T}$ is symmetric and is independent of $\alpha$. Let $\phi_{1},...,\phi_{h},...\in{\mathbb{R}}^{T}$ be the eigenvectors of $Z_{T}$ ordered by the magnitude of their corresponding eigenvalues. Our basis for representing $\mu_{\alpha}$ is spanned by the eigenvectors of $Z_{T}$, and we proceed to study their properties.
Figure 11.1: The filters obtained by the eigenvectors of $Z_{T}$.
@@ 11.1.1 The Magic of Hankel Matrices
The matrix $Z_{T}$ is called the Hilbert matrix, and it is a special case of a Hankel Matrix, which is a square Hermitian matrix with each ascending skew-diagonal from left to right having the same value. To spell out $Z_{T}$ more precisely,
$Z_{T}$ $=\int_{0}^{1}\begin{bmatrix}1&\alpha&\alpha^{2}&&\\ \alpha&\alpha^{2}&\alpha^{3}&&\\ \alpha^{2}&\alpha^{3}&\alpha^{4}&&\\ &&&...&\\ &&&&\alpha^{2T-2}\end{bmatrix}d\alpha=\begin{bmatrix}1&\frac{1}{2}&\frac{1}{3}&&\\ \frac{1}{2}&\frac{1}{3}&\frac{1}{4}&&\\ \frac{1}{3}&\frac{1}{4}&\frac{1}{5}&&\\ &&&...&\\ &&&&\frac{1}{2T-1}\end{bmatrix}$.
@@
Let $\sigma_{k}$ be the $k$-th largest singular value of the Hilbert matrix $Z_{T}\in{\mathbb{R}}^{T\times T}$, then for $k,T>1$:
$\sigma_{k}(Z_{T})\leq 4\pi^{2}\cdot e^{-\frac{k}{\log T}}$
@@ 11.2 Spectral Predictors
Define the class of spectral linear predictors of order $h$ as
@@
Here $\tilde{\mathbf{u}}_{t}^{\top}\phi_{i}$ evaluates the inner product of the $i$’th eigenvector of $Z_{T}$ with the padded history vector $\tilde{\mathbf{u}}_{t}$, where $\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0&\ldots&0\end{bmatrix}^{\top}\in{\mathbb{R}}^{T}$.
Notice that as $h\to T$, this class is equivalent to the class of all linear predictors on the inputs $\Pi^{LIN}$. The reader may wonder at this point: Since a linear transformation composed with another linear transform remains linear, what is the benefit here?
The answer is that linear transformations are equivalent when we have full-rank transforms. In our case, however, we consider low-rank approximations to the transform, and the approximation properties of different subspaces can differ drastically! This is the case in our setting: the spectral basis arising from the matrix $Z_{T}$ gives excellent approximations to the class of (standard) linear predictors, as we formally prove next. The following Theorem is analogous to Lemma 9.4, with the significant difference that the number of parameters $h$ does not depend on the stabilizability parameter of the system $\gamma$ at all!
Theorem 11.3.
Suppose that $\mathbf{x}_{0}=0$ and $|u_{t}|\leq 1$ for all $t$. Then the class $\Pi_{h}$ for $h=O\!\left(\log^{2}\frac{T}{\varepsilon}\right)$ $\varepsilon$-approximates the class $\Pi_{\star}$.
The proof of this lemma relies on the following crucial property of the spectral basis.
@@
Let $\{\phi_{j}\}$ be the orthonormal eigenvectors of $Z_{T}$, with corresponding eigenvalues $\sigma_{j}$. Then for all $j\in[T]$, $j,T>1$, and $\alpha\in[0,1]$, we have
$|\phi_{j}^{\top}\mu_{\alpha}|\;\leq\;(4T^{2}\sigma_{j})^{1/4}.$
In particular, using Theorem 11.1,
$|\phi_{j}^{\top}\mu_{\alpha}|=O\!\left(T^{1/2}e^{-\frac{j}{4\log T}}\right).$
Using this property, we can prove Theorem 11.3 by shifting to a representation on the spectral basis as follows.
Proof of Theorem 11.3.
Consider any sequence and any predictor from $\Pi_{\star}$, say $\pi_{1}=\pi_{A,B,C,0}$. Following equation (11.1), we can write
$\mathbf{y}_{t}^{\pi_{1}}$ $=\beta\cdot\tilde{\mathbf{u}}_{t}^{\top}\mu_{\alpha}$
$=\beta\cdot\tilde{\mathbf{u}}_{t}^{\top}\left(\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top}\right)\mu_{\alpha}\qquad\mbox{(basis property }\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top}=I)$
$=\sum_{i=1}^{T}M_{i}^{\star}\big(\tilde{\mathbf{u}}_{t}^{\top}\phi_{i}\big),\qquad M_{i}^{\star}=\beta(\phi_{i}^{\top}\mu_{\alpha}).$
By Lemma 11.4,
$|M_{i}^{\star}|$ $=|\beta|\,|\phi_{i}^{\top}\mu_{\alpha}|\leq|\beta|\,(4T^{2}\sigma_{i})^{1/4}$
$=O\!\left(|\beta|\,T^{1/2}e^{-\frac{i}{4\log T}}\right).$ (11.2)
Consider the predictor $\pi_{2}\in\Pi_{h}$ that keeps only the first $h$ coefficients. Then we have
$|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}|$ $=\left|\sum_{i=1}^{T}M_{i}^{\star}\big(\tilde{\mathbf{u}}_{t}^{\top}\phi_{i}\big)-\sum_{i=1}^{h}M_{i}^{\star}\big(\tilde{\mathbf{u}}_{t}^{\top}\phi_{i}\big)\right|$
$=\left|\sum_{i=h+1}^{T}M_{i}^{\star}\big(\tilde{\mathbf{u}}_{t}^{\top}\phi_{i}\big)\right|$
$\leq\sum_{i=h+1}^{T}|M_{i}^{\star}|\,\big|\tilde{\mathbf{u}}_{t}^{\top}\phi_{i}\big|$
$\leq\sum_{i=h+1}^{T}|M_{i}^{\star}|\,\|\phi_{i}\|_{2}\|\tilde{\mathbf{u}}_{t}\|_{2}$ (Cauchy–Schwarz)
$\leq\sqrt{T}\,\sum_{i=h+1}^{T}|M_{i}^{\star}|$ $\|\phi_{i}\|_{2}=1,\;|u_{t}|\leq 1$
$\leq\sqrt{T}\,|\beta|\,T^{1/2}\sum_{i=h+1}^{\infty}e^{-\frac{i}{4\log T}}$ (11.2)
The geometric tail can be strictly bounded via an integral:
$\sum_{i=h+1}^{\infty}e^{-\frac{i}{4\log T}}\leq\int_{h}^{\infty}e^{-\frac{x}{4\log T}}dx=4\log T\cdot e^{-\frac{h}{4\log T}}.$
Thus, the overall error is bounded by:
$|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}|\leq|\beta|\,T\big(4\log T\big)e^{-\frac{h}{4\log T}}.$
Choosing
$h\geq 4\log T\left(\log T+\log(4\log T)+\log\frac{|\beta|}{\varepsilon}\right)$
ensures the per-round error is at most $\varepsilon$, completing the proof. ∎
@@
Consider the scalar function
$g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}$
over the interval $[0,1]$.
First,
$\int_{0}^{1}g(\alpha)\,d\alpha$ $=\int_{0}^{1}(\phi_{j}^{\top}\mu_{\alpha})^{2}d\alpha$
$=\phi_{j}^{\intercal}\left(\int_{0}^{1}\mu_{\alpha}\mu_{\alpha}^{\intercal}d\alpha\right)\phi_{j}$
$=\phi_{j}^{\intercal}Z_{T}\phi_{j}=\sigma_{j}.$
Next, we bound the Lipschitz constant of $g$.
We have
$\|\mu_{\alpha}\|_{2}^{2}=\sum_{t=0}^{T-1}\alpha^{2t}\leq T,$
and
$\left\|\frac{\partial}{\partial\alpha}\mu_{\alpha}\right\|_{2}^{2}=\sum_{t=0}^{T-1}(t\alpha^{t-1})^{2}\leq T^{3}.$
Thus
$|g^{\prime}(\alpha)|$ $=2|\mu_{\alpha}^{\top}\phi_{j}|\left|\frac{\partial}{\partial\alpha}(\mu_{\alpha}^{\top}\phi_{j})\right|$
$\leq 2\|\mu_{\alpha}\|_{2}\|\phi_{j}\|_{2}\left\|\frac{\partial}{\partial\alpha}\mu_{\alpha}\right\|_{2}$
$\leq 2\sqrt{T}\sqrt{T^{3}}=2T^{2}.$
Hence $g$ is $L$-Lipschitz with $L\leq 2T^{2}$.
Let $g_{\max}=\max_{\alpha\in[0,1]}g(\alpha)$. Among all nonnegative $L$-Lipschitz functions attaining value $g_{\max}$, the minimal area is achieved by a triangle. Because the half-width of this triangle’s base is at most $g_{\max}/L\leq T/(2T^{2})=1/(2T)\leq 1/2$, the triangle cannot be truncated by both boundaries of the interval $[0,1]$ simultaneously. Thus, the area strictly within $[0,1]$ is at least half the total area of the unbounded triangle, yielding
$\int_{0}^{1}g(\alpha)\,d\alpha\geq\frac{g_{\max}^{2}}{2L}.$
Therefore,
$g_{\max}^{2}\leq 2L\sigma_{j}\leq 4T^{2}\sigma_{j}.$
Taking fourth roots,
$\max_{\alpha}|\phi_{j}^{\top}\mu_{\alpha}|\leq(4T^{2}\sigma_{j})^{1/4}.$
Using Theorem 11.1,
$\sigma_{j}\leq 4\pi^{2}e^{-\frac{j}{\log T}},$
which yields
$|\phi_{j}^{\top}\mu_{\alpha}|=O\!\left(T^{1/2}e^{-\frac{j}{4\log T}}\right).$